# On-the-fly Surrogation for Complex Nonlinear Dynamics*

### v2

@@ V Conclusions
We have presented a novel method that, based on a set of system linearizations, approximates the underlying nonlinear dynamics with a surrogate model that is not pre-computed, i.e., the surrogate model is computed “on-the-fly” when required. Compared to other methods, the main advantages of our approach are: (i) it approximates the global instead of the local underlying nonlinear dynamics, and we show that the proposed surrogate model converges to the true nonlinear system as the number of system linearizations increases; (ii) it avoids computationally intensive processes, such as the training of neural networks; (iii) the inclusion of new system observations only requires solving a linear least squares problem and is therefore straightforward; and (iv) the training data only consists of systems linearizations, which can be efficiently obtained from some simulation environments. In addition, we have demonstrated the global approximation capabilities of our method in various simulation experiments. For further research, we plan to investigate error estimates for the proposed surrogate modelling approach and the combination with state-order reduction techniques to enhance the scalability of our approach.

### v3

@@ V Conclusions
We have presented a novel method that, based on a finite set of system linearizations, which can be efficiently obtained from some simulation environments even if the explicit equations are not available, approximates the underlying nonlinear dynamics with a surrogate model that is computed “on-the-fly” when required. Compared to other methods, the main advantages of our approach are: (i) it approximates the global instead of the local underlying nonlinear dynamics, and we show that the proposed surrogate model converges to the true nonlinear system as the number of system linearizations increases; (ii) it avoids computationally intensive processes, such as the training of neural networks, and (iii) the inclusion of new system observations only requires solving a linear least squares problem and is therefore straightforward. In addition, we have demonstrated the global approximation capabilities of our method in various simulation experiments. For further research, we plan to investigate error estimates for the proposed surrogate modelling approach, the combination with state-order reduction techniques to enhance the scalability of our approach and to introduce sparsity in the interpolant $\mathcal{I}^{\mathcal{D}}$ to reduce the computational cost.